# Distributed games with jumps: an $\alpha$-potential game Approach

### v1

@@ 6. Application to Game-theoretic Motion Planning
$J_{i}({\textbf{{u}}})\coloneqq{\mathbb{E}}\left[\int_{0}^{T}\left(\ell_{i}(u_{i,t})+\frac{1}{N-1}\sum_{j=1,j\neq i}^{N}q_{ij}K(X^{u_{i}}_{i,t}-X^{u_{j}}_{j,t})\right)\,{\mathrm{d}}t+g_{i}(X^{u_{i}}_{i,T})\right],$ (6.1)
@@
${\mathrm{d}}X_{i,t}$ $=b_{i}(t)u_{i,t}{\mathrm{d}}t+\sigma_{i}(t){\mathrm{d}}W_{t}+\sum_{j=1}^{m}\int_{{\mathbb{R}}_{0}^{p}}\gamma_{ij}(t,z)\widetilde{\eta}_{j}({\mathrm{d}}t,{\mathrm{d}}z),\quad t\in(0,T];\quad X_{i,0}=x_{i},$ (6.2)
@@
$\mathcal{A}_{i}=\{u:\Omega\times[0,T]\to A_{i}\mid u\in\mathcal{H}^{2}({\mathbb{R}}^{k}),\|u\|_{\mathcal{H}^{2}({\mathbb{R}}^{k})}\leq U\},$ (6.3)
@@
$K(z)=\exp\left(-\rho{|z|^{2}}\right),\quad\textnormal{with $\rho>0$},$ (6.4)
@@
$K(z)\coloneqq\int_{{\mathbb{R}}^{d}}\mathbf{1}_{B_{r}}(z-v)\gamma_{\delta}(v){\mathrm{d}}v,$ (6.5)
where $\gamma_{\delta}(v)\coloneqq\frac{1}{\delta}\gamma\left(\frac{v}{\delta}\right)$ is a mollifier, with $\gamma:{\mathbb{R}}^{d}\to{\mathbb{R}}$ being a smooth function with compact support, and $\mathbf{1}_{B_{r}}$ is the indicator of the ball $B_{r}$ centered at 0 with radius $r>0$. This kernel function (6.5) has been used in the nonlocal aversion model [3], which captures the phenomenon that each pedestrian is only affected by crowding within their personal space $B_{r}$.
@@ 6.1. Quantifying $\alpha$
For all $i,j\in[N]$, the set $A_{i}$ and the functions $b_{i},\sigma_{i}$ and $\gamma_{ij}$ satisfy (H.11). The functions $(\ell_{i})_{i\in[N]}$, $K$ and $(g_{i})_{i\in[N]}$ are twice continuously differentiable with bounded second-order derivatives.
@@
$\alpha_{N}\leq\frac{1}{2}TB^{2}U^{2}\frac{\kappa}{N-1}\max_{i\in[N]}\sum_{j\neq i}|q_{ji}-q_{ij}|.$ (6.6)
@@
[…]
@@
    $|q_{ij}-q_{ji}|\leq w_{i,j}\frac{1}{c(i,j)^{\beta}},\quad\forall i,j\in{[N]},i\not=j,$ (6.9)
@@
$\alpha_{N}\leq\frac{1}{2}\kappa TB^{2}U^{2}\zeta_{N},$
@@
$\zeta_{N}\leq\frac{C}{N}\max_{i\in[N]}\sum_{j\neq i}\rho^{c(i,j)}\leq\frac{C}{N}\sum_{\ell=1}^{L}\rho^{\ell}d_{G}^{\ell},$ (6.10)
@@
$\zeta_{N}\leq\frac{CL}{N}=\mathcal{O}\left(\frac{\ln N}{N\ln d_{G}}\right).$ (6.11)
@@
$\zeta_{N}\leq\begin{cases}C\frac{1}{d_{G}^{L}}\rho d_{G}\frac{(\rho d_{G})^{L}-1}{\rho d_{G}-1}\leq C\rho^{L}=\mathcal{O}\left(N^{\frac{\ln\rho}{\ln d_{G}}}\right),&\textnormal{if }\rho d_{G}>1,\\ \frac{C}{N}\frac{\rho d_{G}}{1-\rho d_{G}}=\mathcal{O}\left(N^{-1}\right),&\textnormal{if }\rho d_{G}<1.\end{cases}$ (6.12)
@@
$\zeta_{N}\leq\frac{2}{N}\sum_{j\not=i}|q_{ij}-q_{ji}|\leq\frac{2}{N}\left(\max_{i,j\in[N]}|w_{ij}|\right)\sum_{\ell=1}^{N}\frac{n_{\ell}}{\ell^{\beta}}\leq 2\left(\max_{i,j\in[N]}|w_{ij}|\right)\frac{1}{N}\sum_{\ell=1}^{L}\frac{d_{G}^{\ell}}{\ell^{\beta}},$ (6.13)
@@
$\sum_{\ell=1}^{L}\frac{(d_{G})^{\ell}}{\ell^{\beta}}\leq\sum_{\ell=1}^{L-M}\frac{(d_{G})^{\ell}}{\ell^{\beta}}+M\frac{(d_{G})^{L}}{L^{\beta}}$ (6.14)
Since $x\to 1/x^{\beta}$ is deceasing on (0,)(0$\infty$)( 0 , ∞ ), the first term on the right-hand side of (6.14) can be upper bounded by
@@
$\sum_{\ell=1}^{L}\frac{(d_{G})^{\ell}}{\ell^{\beta}}\leq C\left((d_{G})^{L-M^{*}}+\ln L\frac{(d_{G})^{L}}{L^{\beta}}\right)\leq C\ln L\frac{(d_{G})^{L}}{L^{\beta}},$ (6.15)
@@
$\sum_{\ell=1}^{L}\frac{(d_{G})^{\ell}}{\ell^{\beta}}\leq C\left((d_{G})^{L-M^{*}}\ln L+\ln L\frac{(d_{G})^{L}}{L}\right)\leq C\ln L\frac{(d_{G})^{L}}{L},$ (6.16)
@@
$\sum_{\ell=1}^{L}\frac{(d_{G})^{\ell}}{\ell^{\beta}}\leq C\left((d_{G})^{L-M^{*}}L^{1-\beta}+\ln L\frac{(d_{G})^{L}}{L^{\beta}}\right)\leq C\ln L\frac{(d_{G})^{L}}{L^{\beta}},$ (6.17)
@@ 6.2. Numerical results for NEs
${\mathrm{d}}X_{i,t}^{u_{i}}$ $=u_{i,t}{\mathrm{d}}t+\sigma_{i}{\mathrm{d}}W^{i}_{t}+\gamma_{i}{\mathrm{d}}\widetilde{\eta}_{i,t}+\gamma_{0}{\mathrm{d}}\widetilde{\eta}_{0,t},\quad X_{i,0}=x_{i,0},$ (6.18)
@@
Figure 1 illustrates the equilibrium trajectories of the players, where positions at times $t=0.25,0.5,0.75$ are marked by symbols 1, 2, and 3, respectively. The left panel shows the mean positions computed over 500500500 sample trajectories, while the right panel presents a representative single-sample trajectory.
All players begin at the same initial location (indicated by a red circle at position (0,0)(0,0)( 0 , 0 )) and aim to reach a common target (marked by a red cross at (0.5,0.5)(0.5,0.5)( 0.5 , 0.5 )). Early in the game, players disperse in different directions to reduce crowding, a behavior induced by the pairwise aversion term in the cost function. Notably, Player 4 takes a wide detour to avoid other players before converging near the destination. The group exhibits loose coordination: although all players share the same goal, their individual trajectories reflect mutual avoidance dynamics.
Figure 1. Equilibrium trajectories in the aversion game with a Gaussian kernel and uniform interaction weights. Left: mean positions over 500 simulations. Right: one representative trajectory. The solid circle denotes the shared initial position; the cross marks the common target. Markers “1”, “2”, and “3” indicate positions at times 0.250.250.25, 0.50.50.5, and 0.750.750.75, respectively.
@@
The second example considers a flocking game where all players start from the same initial location $x_{i,0}=(0,0)$, and aim for distinct individual target: (0.25,0)(0.25,0)( 0.25 , 0 ), (0,0.5)(0,0.5)( 0 , 0.5 ), (0.5,0)(-0.5,0)( - 0.5 , 0 ), and (0,1)(0,-1)( 0 , - 1 ). Each player is influenced only by idiosyncratic noise, with parameters set as $\sigma_{i}=0.1(i-1)/N$, $\gamma_{i}=0.1$, and $\gamma_{0}=0$ in (6.18). The flocking behavior is modeled using the quadratic kernel $K(z)=\frac{1}{2}|z|^{2}$. Each player $i$ incurs a running cost on control given by $c_{i}(a)=\frac{0.1}{2}|a|^{2}$, and a terminal cost defined by (6.19), with $c_{i}=40$.
We consider two different settings for the interaction weights $(q_{ij})_{i,j=1}^{4}$ in (6.1). In the first setting, uniform interaction is assumed, with $q_{ij}=1$ for all $i\neq j$, so that each player is equally influenced by every other player. In the second setting, a two-group structure is imposed: players 2 and 3 form one group, and players 1 and 4 form another. In this case, $q_{ij}=1$ if players $i$ and $j$ belong to the same group, and $q_{ij}=0$ otherwise. This models selective flocking behavior, where players tend to coordinate only with those in their own group.
@@
Figure 4 presents two sample trajectories of the resulting equilibrium dynamics. The common jumps introduce abrupt, synchronized shifts in the players’ positions, followed by realignment as they continue moving toward their respective targets. While the jump events cause irregularities in the intermediate paths, the overall flocking behavior remains consistent with the patterns observed in the present setting with pure idiosyncratic noises.

### v2

@@ 6. Application to Game-theoretic Motion Planning
(6.1) $J_{i}({\textbf{{u}}})\coloneqq{\mathbb{E}}\left[\int_{0}^{T}\left(\ell_{i}(u_{i,t})+\frac{1}{N-1}\sum_{j=1,j\neq i}^{N}q_{ij}K(X^{u_{i}}_{i,t}-X^{u_{j}}_{j,t})\right)\,{\mathrm{d}}t+g_{i}(X^{u_{i}}_{i,T})\right],$
@@
(6.2) ${\mathrm{d}}X_{i,t}$ $=b_{i}(t)u_{i,t}{\mathrm{d}}t+\sigma_{i}(t){\mathrm{d}}W_{t}+\sum_{j=1}^{m}\int_{{\mathbb{R}}_{0}^{p}}\gamma_{ij}(t,z)\widetilde{\eta}_{j}({\mathrm{d}}t,{\mathrm{d}}z),\quad t\in(0,T];\quad X_{i,0}=x_{i},$
@@
(6.3) $\mathcal{A}_{i}=\{u:\Omega\times[0,T]\to A_{i}\mid u\in\mathcal{H}^{2}({\mathbb{R}}^{k}),\|u\|_{\mathcal{H}^{2}({\mathbb{R}}^{k})}\leq U\},$
@@
(6.4) $K(z)=\exp\left(-\rho{|z|^{2}}\right),\quad\textnormal{with $\rho>0$},$
@@
(6.5) $K(z)\coloneqq\int_{{\mathbb{R}}^{d}}\mathbf{1}_{B_{r}}(z-v)\gamma_{\delta}(v){\mathrm{d}}v,$
where $\gamma_{\delta}(v)\coloneqq\frac{1}{\delta}\gamma\left(\frac{v}{\delta}\right)$ is a radially symmetric mollifier, with $\gamma:{\mathbb{R}}^{d}\to{\mathbb{R}}$ being a smooth function with compact support, and $\mathbf{1}_{B_{r}}$ is the indicator of the ball $B_{r}$ centered at 0 with radius $r>0$. This kernel function (6.5) has been used in the nonlocal aversion model [3], which captures the phenomenon that each pedestrian is only affected by crowding within their personal space $B_{r}$.
@@ 6.1. Quantifying $\alpha$
For all $i,j\in[N]$, the set $A_{i}$ and the functions $b_{i},\sigma_{i}$ and $\gamma_{ij}$ satisfy (H.11). The functions $(\ell_{i})_{i\in[N]}$, $K$ and $(g_{i})_{i\in[N]}$ are twice continuously differentiable with bounded second-order derivatives, and K(z)=K(z)K(z)=K(-z) for all $z\in{\mathbb{R}}^{d}$.
@@
(6.6) $\alpha_{N}\leq TB^{2}U^{2}\frac{\kappa}{N-1}\max_{i\in[N]}\sum_{j\neq i}|q_{ji}-q_{ij}|.$
@@
    Asymmetric interaction. To capture asymmetric interactions, we assume that the interaction weights are determined by an underlying undirected graph $G$, where the vertices represent the set of players $[N]$, and each edge indicates a connectivity relation between the corresponding players.
@@
    (6.9) $|q_{ij}-q_{ji}|\leq w_{i,j}\frac{1}{c(i,j)^{\beta}},\quad\forall i,j\in{[N]},i\not=j,$
@@
$\alpha_{N}\leq\kappa TB^{2}U^{2}\zeta_{N},$
@@
(6.10) $\zeta_{N}\leq\frac{C}{N}\max_{i\in[N]}\sum_{j\neq i}\rho^{c(i,j)}\leq\frac{C}{N}\sum_{\ell=1}^{L}\rho^{\ell}d_{G}^{\ell},$
@@
(6.11) $\zeta_{N}\leq\frac{CL}{N}=\mathcal{O}\left(\frac{\ln N}{N\ln d_{G}}\right).$
@@
(6.12) $\zeta_{N}\leq\begin{cases}C\frac{1}{d_{G}^{L}}\rho d_{G}\frac{(\rho d_{G})^{L}-1}{\rho d_{G}-1}\leq C\rho^{L}=\mathcal{O}\left(N^{\frac{\ln\rho}{\ln d_{G}}}\right),&\textnormal{if }\rho d_{G}>1,\\ \frac{C}{N}\frac{\rho d_{G}}{1-\rho d_{G}}=\mathcal{O}\left(N^{-1}\right),&\textnormal{if }\rho d_{G}<1.\end{cases}$
@@
(6.13) $\zeta_{N}\leq\frac{2}{N}\sum_{j\not=i}|q_{ij}-q_{ji}|\leq\frac{2}{N}\left(\max_{i,j\in[N]}|w_{ij}|\right)\sum_{\ell=1}^{N}\frac{n_{\ell}}{\ell^{\beta}}\leq 2\left(\max_{i,j\in[N]}|w_{ij}|\right)\frac{1}{N}\sum_{\ell=1}^{L}\frac{d_{G}^{\ell}}{\ell^{\beta}},$
@@
(6.14) $\sum_{\ell=1}^{L}\frac{(d_{G})^{\ell}}{\ell^{\beta}}\leq\sum_{\ell=1}^{L-M}\frac{(d_{G})^{\ell}}{\ell^{\beta}}+M\frac{(d_{G})^{L}}{L^{\beta}}$
Since $x\to 1/x^{\beta}$ is deceasing on $(0,\infty)$, the first term on the right-hand side of (6.14) can be upper bounded by
@@
(6.15) $\sum_{\ell=1}^{L}\frac{(d_{G})^{\ell}}{\ell^{\beta}}\leq C\left((d_{G})^{L-M^{*}}+\ln L\frac{(d_{G})^{L}}{L^{\beta}}\right)\leq C\ln L\frac{(d_{G})^{L}}{L^{\beta}},$
@@
(6.16) $\sum_{\ell=1}^{L}\frac{(d_{G})^{\ell}}{\ell^{\beta}}\leq C\left((d_{G})^{L-M^{*}}\ln L+\ln L\frac{(d_{G})^{L}}{L}\right)\leq C\ln L\frac{(d_{G})^{L}}{L},$
@@
(6.17) $\sum_{\ell=1}^{L}\frac{(d_{G})^{\ell}}{\ell^{\beta}}\leq C\left((d_{G})^{L-M^{*}}L^{1-\beta}+\ln L\frac{(d_{G})^{L}}{L^{\beta}}\right)\leq C\ln L\frac{(d_{G})^{L}}{L^{\beta}},$
@@ 6.2. Numerical results for NEs
(6.18) ${\mathrm{d}}X_{i,t}^{u_{i}}$ $=u_{i,t}{\mathrm{d}}t+\sigma_{i}{\mathrm{d}}W^{i}_{t}+\gamma_{i}{\mathrm{d}}\widetilde{\eta}_{i,t}+\gamma_{0}{\mathrm{d}}\widetilde{\eta}_{0,t},\quad X_{i,0}=x_{i,0},$
@@
Figure 1 illustrates the equilibrium trajectories of the players, where positions at times $t=0.25,0.5,0.75$ are marked by symbols 1, 2, and 3, respectively. The left panel shows the mean positions computed over $500$ sample trajectories, while the right panel presents a representative single-sample trajectory.
All players begin at the same initial location (indicated by a red circle at position $(0,0)$) and aim to reach a common target (marked by a red cross at $(0.5,0.5)$). Early in the game, players disperse in different directions to reduce crowding, a behavior induced by the pairwise aversion term in the cost function. Notably, Player 4 takes a wide detour to avoid other players before converging near the destination. The group exhibits loose coordination: although all players share the same goal, their individual trajectories reflect mutual avoidance dynamics.
Figure 1. Equilibrium trajectories in the aversion game with a Gaussian kernel and uniform interaction weights. Left: mean positions over 500 simulations. Right: one representative trajectory. The solid circle denotes the shared initial position; the cross marks the common target. Markers “1”, “2”, and “3” indicate positions at times $0.25$, $0.5$, and $0.75$, respectively.
@@
The second example considers a flocking game where all players start from the same initial location $x_{i,0}=(0,0)$, and aim for distinct individual target: $(0.25,0)$, $(0,0.5)$, (0.5,0)(-0.5,0), and (0,1)(0,-1). Each player is influenced only by idiosyncratic noise, with parameters set as $\sigma_{i}=0.1(i-1)/N$, $\gamma_{i}=0.1$, and $\gamma_{0}=0$ in (6.18). The flocking behavior is modeled using the quadratic kernel $K(z)=\frac{1}{2}|z|^{2}$. Each player $i$ incurs a running cost on control given by $\ell_{i}(a)=\frac{0.1}{2}|a|^{2}$, and a terminal cost defined by (6.19), with $c_{i}=40$.
We consider two different settings for the interaction weights $(q_{ij})_{i,j=1}^{N}$ in (6.1). In the first setting, uniform interaction is assumed, with $q_{ij}=1$ for all $i\neq j$, so that each player is equally influenced by every other player. In the second setting, a two-group structure is imposed: players 2 and 3 form one group, and players 1 and 4 form another. In this case, $q_{ij}=1$ if players $i$ and $j$ belong to the same group, and $q_{ij}=0$ otherwise. This models selective flocking behavior, where players tend to coordinate only with those in their own group.
@@
Figure 4 presents two sample trajectories of the resulting equilibrium dynamics. The common jumps introduce abrupt, synchronized shifts in the players’ positions, followed by realignment as they continue moving toward their respective targets. While the jump events cause irregularities in the intermediate paths, the overall flocking behavior remains consistent with the patterns observed in Figure 2 for the setting with purely idiosyncratic noise.